# Weak-Coupling Limit of the Lattice Nonlinear Schrödinger Integral Equation

### 10 Mar 2026 v1

@@ 6.3 Numerical verification
Table 1: Effective constant $C_{\mathrm{eff}}(Q):=\tilde{\rho}(0;Q)-(\log Q)/\pi$ and its deviation from $C_{*}=(\gamma_{E}+\log 2)/\pi$.
@@
Table 2: Three-point Richardson extrapolation for $C$. The best result ($Q=80,150,300$) gives $C_{\mathrm{extr}}=0.404\,369\,014$, agreeing with $(\gamma_{E}+\log 2)/\pi=0.404\,369\,053$ to 8 significant figures.
Fixing $C=C_{*}$ and fitting the correction parameters by least squares gives residuals with maximum absolute value $2.8\times 10^{-6}$, confirming the prediction to the limits of double-precision arithmetic. For contrast, fixing $C=\gamma_{E}/\pi$ (without the $\log 2$ term) yields a maximum residual of $1.5\times 10^{-1}$—five orders of magnitude worse, ruling out $C=\gamma_{E}/\pi$ decisively.
@@ 7 Energy and physical predictions
Having determined the constant $C$ in the peak density expansion (75), we now derive the ground-state energy and discuss the physical consequences. The central tool is an exact identity relating the energy integral to the peak density, which holds for all $Q>0$ and reduces the energy calculation to a single quantity already determined in Section 6.
@@ 7.1 Energy identity
In the rescaled variables (23), the physical energy per site (22) takes the form $e(\kappa)=-E_{\mathrm{inner}}/\kappa$, where
$E_{\mathrm{inner}}(Q):=\int_{-Q}^{Q}\frac{2}{1+\xi^{2}}\,\tilde{\rho}(\xi;\,Q)\,d\xi$ (101)
is the rescaled energy integral. Note that the integrand involves the Lorentzian weight (26), a fact that leads to a remarkable simplification.
@@
where the last step uses the definition (101). Rearranging gives the exact, non-asymptotic identity
@@
valid for all $Q>0$. This identity reduces the energy calculation to the peak density $\tilde{\rho}(0;\,Q)$—determined in Section 6.
@@
7.2 Asymptotic expansion and mode-sum derivation
Substituting the peak density (75) with (76) into the identity (103):
@@ 7.2 Asymptotic expansion and mode-sum derivation
which gives $E_{\mathrm{inner}}=2\log(2Q)+2\gamma_{E}-2+\mathcal{O}(1/Q)$, confirming (104).
7.3 Ground-state energy
@@
$\rho(0;\,\kappa)=\frac{\log(2q/\kappa)+\gamma_{E}}{\pi\kappa}+\cdots\,.$ (107)
The physical ground-state energy per site is $e(\kappa)=-E_{\mathrm{inner}}/\kappa$:
$e(\kappa)=-\frac{E_{\mathrm{inner}}}{\kappa}=-\frac{2}{\kappa}\left[\log(2q/\kappa)+\gamma_{E}-1\right]+\cdots\,.$ (108)
@@ 7.4 Energy at fixed particle density and comparison with Lieb–Liniger
The physically natural regime is fixed particle density $D_{0}$, not fixed $q$. Substituting the relation $q(\kappa)=D_{0}\kappa-(\kappa/2\pi)\log(1/\kappa)+\cdots$ into (108) gives, to leading order,
$e(\kappa)=-\frac{2}{\kappa}\left[\log(2D_{0})+\gamma_{E}-1\right]+\cdots\,$, (109)
The bracketed factor $\log(2D_{0})+\gamma_{E}-1]$ is independent of $\kappa$, so the energy diverges as $1/\kappa$ at fixed density—reflecting the fact that the lattice model at small $\kappa$ accommodates $O(1/\kappa)$ particles per site, while the energy per particle remains $O(1)$. This should be contrasted with the Lieb–Liniger energy $e\sim\gamma=c/D_{0}$ at weak coupling, which vanishes linearly. The subleading correction in (109) arises from the logarithmic term in $q(\kappa)$; its detailed evaluation requires the constant $b$ in the density expansion (67), which is not yet known analytically.
These results are qualitatively different from the continuous Lieb–Liniger expansion at every level, as summarised in Table 3.
@@ 8.1 Instanton action from the Wiener–Hopf symbol
The edge boundary layer at $\xi=\pm Q$ is governed by the Wiener–Hopf symbol $\Sigma(p)=1-e^{-|p|}$. On the real axis, $\Sigma(p)=0$ only at $p=0$. However, the analytic continuation to the complex $p$-plane has additional zeros at $p_{n}=2\pi in$ for $n\in\mathbb{Z}\setminus\{0\}$. The nearest zeros to the real axis lie at $p_{\pm 1}=\pm 2\pi i$, at distance $2\pi$ from the real axis.
@@
The same value was found in [4] for the non-perturbative corrections to the circular disc capacitor, whose integral equation is the Love equation—dual to our lattice NLS equation. This coincidence is not accidental: the two problems share the same Wiener–Hopf symbol.
8.2 Spectral gap: algebraic closure with exponential correction
The largest eigenvalue $\lambda_{0}(Q)$ of $\mathcal{K}_{Q}$ approaches $2\pi$ as $Q\to\infty$, with spectral gap $\Delta_{0}(Q)=2\pi-\lambda_{0}(Q)$. The perturbative (algebraic) component must be distinguished from the non-perturbative (exponential) correction:
@@ 8.2 Spectral gap: algebraic closure with exponential correction
Numerical computation (Figure 2b) shows that $Q\cdot\Delta_{0}$ is not constant, indicating logarithmic corrections to the $1/Q$ law. The ratio $\Delta_{1}/\Delta_{0}$ converges to $\approx 2.37$ as $Q$ grows (computed from the eigenvalue data in Table 4 and confirmed at larger $Q$ values in Figure 2a), confirming that the leading eigenvalue separates from the rest of the spectrum. The exponential correction is therefore a correction to the algebraically closing gap, not its leading behaviour—the same structure as in the disc capacitor problem [4, 37].
8.3 The Fredholm determinant and Szegő asymptotics
@@
provides a global characterisation of the operator spectrum. The first Szegő limit theorem [80, 13] gives
$\frac{1}{2Q}\log\mathcal{F}(Q)\;\to\;\frac{1}{\pi}\int_{0}^{\infty}\log(1-e^{-p})\,dp=-\frac{\pi}{6}\,$, (113)
using $\int_{0}^{\infty}\log(1-e^{-p})\,dp=-\zeta(2)=-\pi^{2}/6$.
@@
where $\sigma$ is the transseries parameter and the $k=0$ sector reproduces the perturbative expansion. The perturbative sector ($k=0$) and the instanton action $A=2\pi$ are established analytically in this paper; the multi-instanton structure for $k\geq 1$ and the conjectured factorial growth of the perturbative coefficients are based on the structural analogy with the Lieb–Liniger and disc capacitor problems [65, 4]. If the perturbative coefficients grow factorially, $a_{n0}\sim n!\,S_{1}/(2\pi)^{n}$ as $n\to\infty$, then the Borel transform has a singularity at $t=A=2\pi$, rendering the series non-Borel summable along the positive real axis. The imaginary ambiguity of the lateral Borel resummation is then cancelled by the one-instanton contribution, as required by the reality of $\tilde{\rho}(0;\,Q)$.
8.4 Analytical support and connection to the Lieb–Liniger resurgence
In the continuous Lieb–Liniger model [66], the instanton action equals $4\pi$ (in their normalisation) and the Borel-plane singularity is associated with the energy gap of the dual Gaudin–Yang model. The Lieb–Liniger kernel with constant driving term has Wiener–Hopf symbol $\sigma_{\mathrm{LL}}(p)=1-e^{-\kappa|p|}$, which has zeros at $p=2\pi in/\kappa$. At $\kappa=1$ (the rescaled lattice equation), $\sigma(p)=1-e^{-|p|}$ has zeros at $p=2\pi in$—exactly our symbol. Thus the instanton action of the lattice NLS equation at $\kappa=1$ coincides with that of the Lieb–Liniger equation at $\kappa=1$.
By analogy with the Lieb–Liniger / Gaudin–Yang duality, one expects $A=2\pi$ to be related to a spectral gap in a dual model obtained by flipping the sign of the interaction (attractive lattice NLS, or equivalently XXXs=+1).
The analytical prediction $A=2\pi$ rests on solid structural ground—it follows from the zero structure of $\Sigma(p)=1-e^{-|p|}$, is consistent with the disc capacitor results [4], and matches the Lieb–Liniger instanton at $\kappa=1$. Preliminary numerical tests, presented in Appendix D, extract $\sim 5$ stable perturbative coefficients whose alternating signs are consistent with a Borel singularity on the positive real axis, but a definitive numerical confirmation of the instanton action requires extended-precision arithmetic ($\gtrsim 50$ digits) to extract $\sim 20$$30$ perturbative coefficients and detect the instanton directly.
@@ 9 Conclusions
Inner region and the Bose–Einstein distribution. Rescaling via (23), the integral equation becomes a Lieb–Liniger-type equation on [Q,Q][-Q,Q] with $\kappa=1$. On the full line ($Q=\infty$), the Fourier transform of the rescaled solution is the Bose–Einstein distribution (34), whose $1/|p|$ singularity at the origin produces a logarithmic divergence in position space. On the finite domain, this divergence is regulated by the interval length, yielding the asymptotic expansion (75) with the constant $C$ given by (76). This was established by two independent routes: a heuristic mode-counting argument based on (85) combined with a digamma representation of the full-line solution, and an analytical Wiener–Hopf derivation based on the spectral response function of the truncated operator and the normalisation $G_{\pm}(0)=1$ (92) of the regularised Wiener–Hopf factors. Three-point Richardson extrapolation confirms $C$ to eight significant figures.
Outer region and density decomposition. In the outer region $1\ll|\xi|\ll Q$, the solution reduces to a uniform Fermi sea with bulk value (44). Using the duality (53) with the Love integral equation for the circular disc capacitor, together with the self-consistency identity (54) and the integral identity (56), we derived the total density expansion (55), with $b\approx-0.2173$ determined numerically. In physical variables, at fixed Fermi boundary $q$, the density is given by (67).
Edge boundary layer. Near the Fermi boundary $\xi=\pm Q$, the solution drops from its bulk value to zero in a boundary layer of width $\mathcal{O}(1)$ in the rescaled variable. The transition is governed by a half-line Wiener–Hopf problem with symbol (70), whose factorisation (71)–(72) we obtained in terms of gamma functions and logarithmic exponentials. This factorisation controls the mapping between the spatial cutoff $Q$ and the effective harmonic cutoff, and identifies the factor $A_{\mathrm{WH}}=2$ responsible for the $\log 2$ contribution to $C$.
@@
Towards resurgence. The Wiener–Hopf data predicts the instanton action (110), consistent with the disc capacitor results [4] and the Lieb–Liniger resurgence [65]. Preliminary numerical extraction of perturbative coefficients (Appendix D) yields $\sim 5$ stable coefficients with alternating signs consistent with a Borel singularity, but the number of reliably extracted coefficients is presently too small for a definitive confirmation of factorial divergence. Obtaining higher-order coefficients—either through extended-precision numerics or through the explicit Wiener–Hopf solution of the edge problem—and establishing the connection between the instanton action and the spectral gap of a dual model remain important open problems. The analytical determination of the constant $b\approx-0.2173$ in the density expansion—which we conjecture involves the Fisher–Hartwig constant $\beta_{\mathrm{FH}}$ of the Fredholm determinant and possibly the Barnes $G$-function—the closed-form evaluation of the subleading coefficients in the expansion of $\tilde{\rho}(0;Q)$, and the extension of the present analysis to excited states and finite temperature are further natural directions.
@@ A.1 Basic spectral properties
Self-adjointness follows from the symmetry $K(\xi-\eta)=K(\eta-\xi)$; the Hilbert–Schmidt norm is bounded by $\|\mathcal{K}_{Q}\|_{\mathrm{HS}}^{2}\leq 4\pi Q$ since $\int_{-\infty}^{\infty}(1+x^{2})^{-2}\,dx=\pi/2$; positivity follows from the Fourier representation $\langle f,\mathcal{K}_{Q}f\rangle=\int 2\pi\,e^{-|p|}\,|\hat{\tilde{f}}(p)|^{2}\,dp\geq 0$, where $\tilde{f}$ is the extension by zero to $\mathbb{R}$.
@@
and setting $\xi=\eta$ and integrating gives the trace identity $\sum_{n=0}^{\infty}\lambda_{n}(Q)=4Q$.
@@
For a truncated convolution operator on [Q,Q][-Q,Q] with symbol $\widehat{K}(p)=2\pi\,e^{-|p|}$, the first Szegő limit theorem (see [80, 13]) gives the asymptotic distribution of eigenvalues. Define the symbol of the Love operator $\mathcal{L}_{Q}=(2\pi)^{-1}\mathcal{K}_{Q}$ as $\sigma(p)=1-e^{-|p|}$, so the eigenvalues of $\mathcal{L}_{Q}$ are $\mu_{n}=\lambda_{n}/(2\pi)\in(0,1)$.
@@
In the large-$Q$ limit, the operator is “locally equivalent” to $\mathcal{L}$, and for a self-adjoint operator whose symbol takes values in $[0,1]$, the eigenvalue distribution converges (in the sense of (121)) to the push-forward of the Lebesgue measure on momentum space under the symbol map $p\mapsto e^{-|p|}$. The normalisation factor $1/(2Q)$ is the reciprocal of the interval length, reflecting the density of states per unit length.
@@
since $\log(1/(1-\varepsilon))\sim\varepsilon$ for small $\varepsilon$. This confirms that $O(Q)$ eigenvalues accumulate at $2\pi$ for large $Q$.
@@
where $\alpha=2[1-\log 2]$. (One evaluates the integral using $\int\arctan(x+R)\,dx=(x+R)\arctan(x+R)-\tfrac{1}{2}\log(1+(x+R)^{2})+C$ and expands for large $R$.)
@@
The uncertainty principle for functions supported on [Q,Q][-Q,Q] gives $\int|p|^{2}\,|\hat{\tilde{\phi}}_{0}|^{2}\,dp\geq c/Q^{2}$ (since $\phi_{0}$ cannot simultaneously be concentrated in both position and frequency). Combined with (130) via Cauchy–Schwarz, this yields $\delta\geq c_{1}/(Q\,(\log Q)^{2})$ after optimising the frequency partition. ∎
@@
with $c_{0}\approx 6.43$ and $c_{1}\approx 0.15$, rather than a pure power law $c/Q^{\alpha}$. An effective exponent $\alpha>1$ may be observed over limited ranges of $Q$ (particularly at small $Q$, where pre-asymptotic effects are significant), but for large $Q$ the data is consistent with the algebraic rate $1/Q$ modulated by a slowly growing logarithmic correction. This logarithmic correction originates from the interplay between the linear zero of the symbol $\sigma(p)=1-e^{-|p|}$ at $p=0$ and the Fisher–Hartwig correction to the Szegő asymptotics (see Section A.3).

### 22 Mar 2026 v2

@@ 6.3 Numerical verification
Table 1: Effective constant $C_{\mathrm{eff}}(Q):=\tilde{\rho}(0;Q)-(\log Q)/\pi$ and its deviation from $C_{*}=(\gamma_{E}+\log 2)/\pi$ (88).
@@
Table 2: Three-point Richardson extrapolation for $C$. The best result ($Q=80,150,300$) gives $C_{\mathrm{extr}}=0.404\,369\,014$, agreeing with $(\gamma_{E}+\log 2)/\pi=0.404\,369\,053$ to eight significant digits.
Fixing $C=C_{*}$ and fitting the correction parameters by least squares gives residuals with maximum absolute value $2.8\times 10^{-6}$, confirming the prediction to the limits of double-precision arithmetic.
@@ 7 Energy and physical predictions
Having determined the constant $C$ in the peak density expansion (87), we now derive the ground-state energy and discuss the physical consequences. The central tool is the identity relating the energy integral to the peak density, which holds for all $Q>0$ and reduces the energy calculation to a single quantity already determined in Section 6.
@@ 7.1 Energy identity
In the rescaled variables (24), the physical energy per site (23) takes the form
$e(\kappa)=-\frac{E_{\rm inner}}{\kappa}\,,\qquad E_{\mathrm{inner}}(Q):=\int_{-Q}^{Q}\frac{2}{1+\xi^{2}}\,\tilde{\rho}(\xi;\,Q)\,d\xi$ (122)
where $E_{\mathrm{inner}}(Q)$ is the rescaled energy integral. Note that the integrand involves the Lorentzian weight (28), a fact that leads to a remarkable simplification.
@@
where the last step uses the definition (122). Therefore, we have the identity
@@
valid for all $Q>0$. This identity reduces the energy calculation to the peak density $\tilde{\rho}(0;\,Q)$.
@@
Substituting the peak density (87) with ($8$) into the identity (124):
@@
which, for large $Q$ confirms (125).
@@
$\rho(0;\,\kappa)=\frac{1}{\pi\kappa}\left[\log\left(\frac{2q}{\kappa}\right)+\gamma_{E}\right]\,.$ (128)
The physical ground-state energy per site is:
$e(\kappa)=-\frac{E_{\mathrm{inner}}}{\kappa}=-\frac{2}{\kappa}\left[\log\left(\frac{2q}{\kappa}\right)+\gamma_{E}-1\right]\,.$ (129)
@@ 7.2 Energy at fixed particle density and comparison with Lieb–Liniger
The physically natural regime is fixed particle density $D_{0}$, not fixed $q$. Thus, expanding
$q(\kappa)=D_{0}\kappa-\frac{\kappa}{2\pi}\log\left(\frac{1}{\kappa}\right)+\cdots$ (130)
and substituting into (129) yields, to leading order,
$e(\kappa)=-\frac{2}{\kappa}\left(\log(2D_{0})+\gamma_{E}-1\right)\,.$ (131)
Note that the bracketed factor in (131) is independent of $\kappa$, so the energy diverges as $1/\kappa$ at fixed density—reflecting the fact that the lattice model at small $\kappa$ accommodates $\mathcal{O}(1/\kappa)$ particles per site, while the energy per particle remains $\mathcal{O}(1)$. This should be contrasted with the Lieb–Liniger energy $e\sim\gamma=c/D_{0}$ at weak coupling, which vanishes linearly. The subleading correction in (131) arises from the logarithmic term in $q(\kappa)$; its detailed evaluation requires the constant $b$ in the density expansion (75), which is not yet known analytically. These results are qualitatively different from the continuous Lieb–Liniger expansion at every level, as summarised in Table 3.
@@ 8.1 Instanton action from the Wiener–Hopf symbol
The edge boundary layer at $\xi=\pm Q$ is governed by the Wiener–Hopf symbol $\Sigma(p)$ (79). On the real axis, $\Sigma(p)=0$ only at $p=0$. However, the analytic continuation to the complex $p$-plane has additional zeros at $p_{n}=2\pi in$ for $n\in\mathbb{Z}\setminus\{0\}$. The nearest zeros to the real axis lie at $p_{\pm 1}=\pm 2\pi i$, at distance $2\pi$ from the real axis.
@@
The same value was found in [4] for the non-perturbative corrections to the circular disc capacitor, whose integral equation is the Love equation—dual to our lattice NLS equation. This coincidence is not accidental: the two problems share the same Wiener–Hopf symbol. The same instanton action governs the exponential correction to the spectral gap $\Delta_{0}(Q)=2\pi-\lambda_{0}(Q)$: the perturbative (algebraic) component must be distinguished from the non-perturbative (exponential) correction,
@@
Since the algebraic gap closes as $\Delta_{0}^{(\mathrm{pert})}\sim\pi/(Q\log Q)$ (see Section 3.3 and Appendix A), the exponential correction is subleading—the same structure as in the disc capacitor problem [4, 37].
8.2 Fredholm determinant and Szegő asymptotics
@@
provides a global characterisation of the operator spectrum. The first Szegő limit theorem [80, 13] informs us that
$\frac{1}{2Q}\log\mathcal{F}(Q)\;\to\;\frac{1}{\pi}\int_{0}^{\infty}\log(1-e^{-p})\,dp=-\frac{\zeta(2)}{\pi}=-\frac{\pi}{6}\,.$ (135)
@@
where $\sigma$ is the transseries parameter and the $k=0$ sector reproduces the perturbative expansion. The perturbative sector ($k=0$) and the instanton action (132) are established analytically; the multi-instanton structure for $k\geq 1$ and the conjectured factorial growth of the perturbative coefficients are based on the structural analogy with the Lieb–Liniger and disc capacitor problems [65, 4]. If the perturbative coefficients grow factorially, $a_{n0}\sim n!\,S_{1}/(2\pi)^{n}$ as $n\to\infty$, then the Borel transform has a singularity at $t=A=2\pi$, rendering the series non-Borel summable along the positive real axis. The imaginary ambiguity of the lateral Borel resummation is then cancelled by the one-instanton contribution, as required by the reality of $\tilde{\rho}(0;\,Q)$.
In the continuous Lieb–Liniger model [66], the instanton action equals $4\pi$ and the Borel-plane singularity is associated with the energy gap of the dual Gaudin–Yang model. The Lieb–Liniger kernel with constant driving term has Wiener–Hopf symbol $\sigma_{\mathrm{LL}}(p)=1-e^{-\kappa|p|}$, which has zeros at $p=2\pi in/\kappa$. At $\kappa=1$ (the rescaled lattice equation), $\sigma(p)$ matches our symbol (79) and has zeros at $p=2\pi in$. Thus the instanton action of the lattice NLS equation at $\kappa=1$ coincides with that of the Lieb–Liniger equation at $\kappa=1$. By analogy with the Lieb–Liniger / Gaudin–Yang duality, one expects (132) to be related to a spectral gap in a dual model obtained by flipping the sign of the interaction (attractive lattice NLS, or equivalently XXXs=+1).
In summary, the analytical prediction (132) rests on solid structural ground—it follows from the zero structure of $\Sigma(p)$, is consistent with the disc capacitor results [4], and matches the Lieb–Liniger instanton at $\kappa=1$. Preliminary numerical tests, presented in Appendix D, extract $\sim 5$ stable perturbative coefficients whose alternating signs are consistent with a Borel singularity on the positive real axis, but a definitive numerical confirmation of the instanton action requires extended-precision arithmetic ($\gtrsim 50$ digits) to extract $\sim 20$$30$ perturbative coefficients and detect the instanton directly.
@@ 9 Conclusions
Inner region and the Bose–Einstein distribution. Rescaling via (24), the integral equation becomes a Lieb–Liniger-type equation on [Q,Q][-Q,Q] with $\kappa=1$. On the full line ($Q=\infty$), the Fourier transform of the rescaled solution is the Bose–Einstein distribution (36), whose $1/|p|$ singularity at the origin produces a logarithmic divergence in position space. On the finite domain, this divergence is regulated by the interval length, yielding the asymptotic expansion (87) with the constant $C$ given by (88). This was established by two independent routes: a heuristic mode-counting argument based on (100) combined with a digamma representation of the full-line solution, and an analytical Wiener–Hopf derivation based on the spectral response function of the truncated operator and the normalisation $G_{\pm}(0)=1$ (110) of the regularised Wiener–Hopf factors.
Outer region and density decomposition. In the outer region $1\ll|\xi|\ll Q$, the solution reduces to a uniform Fermi sea with bulk value (46). Using the duality (55) with the Love integral equation for the circular disc capacitor, together with the self-consistency identity (56) and the integral identity (58), we derived the total density expansion (57), with $b\approx-0.2173$ determined numerically. The profile integral (72) shows that the inner and outer regions contribute exactly $1/2$ to D(Q)QD(Q)-Q, localising the $(2\pi)^{-1}\log Q$ term and the constant $b$ in the edge boundary layers. In physical variables, at fixed Fermi boundary $q$, the density is given by (75).
Edge boundary layer. Near the Fermi boundary $\xi=\pm Q$, the solution drops from its bulk value to zero in a boundary layer of width $\mathcal{O}(1)$ in the rescaled variable. The transition is governed by a half-line Wiener–Hopf problem with symbol (79), whose factorisation (81)–(82) we obtained. This factorisation controls the mapping between the spatial cutoff $Q$ and the effective harmonic cutoff, and identifies the factor $A_{\mathrm{WH}}=2$ responsible for the $\log 2$ contribution to $C$.
@@
Towards resurgence. The Wiener–Hopf data predicts the instanton action (132), consistent with the disc capacitor results [4] and the Lieb–Liniger resurgence [65]. Preliminary numerical extraction of perturbative coefficients (Appendix D) yields five stable coefficients with alternating signs consistent with a Borel singularity, but the number of reliably extracted coefficients is presently too small for a definitive confirmation of factorial divergence. Obtaining higher-order coefficients—either through extended-precision numerics or through the explicit Wiener–Hopf solution of the edge problem—and establishing the connection between the instanton action and the spectral gap of a dual model remain important open problems. The analytical determination of the constant $b\approx-0.2173$ in the density expansion remains open; the profile integral (72) shows that the inner and outer regions contribute $1/2$ to D(Q)QD(Q)-Q, localising the entire non-trivial content of $b$ in the edge boundary layers and the Fisher–Hartwig constant $\beta_{\mathrm{FH}}$ of the Fredholm determinant. The closed-form evaluation of the subleading coefficients in the expansion of $\tilde{\rho}(0;Q)$, and the extension of the present analysis to excited states and finite temperature are further natural directions.
@@
Data availability
All numerical computations and data reported in this paper are available at github.com/ftahas/lattice-NLS.
@@ A.1 Basic spectral properties
Self-adjointness follows from the symmetry $K(\xi-\eta)=K(\eta-\xi)$. The Hilbert–Schmidt norm is bounded by $\|\mathcal{K}_{Q}\|_{\mathrm{HS}}^{2}\leq 4\pi Q$ since
$\int_{-\infty}^{\infty}\frac{1}{\left(1+x^{2}\right)^{2}}\,dx=\frac{\pi}{2}\,.$ (140)
Positivity follows from the Fourier representation
$\langle f,\mathcal{K}_{Q}f\rangle=\int_{-\infty}^{\infty}2\pi\,e^{-|p|}\,|\hat{\tilde{f}}(p)|^{2}\,dp\geq 0\,$, (141)
where $\tilde{f}$ is the extension by zero to $\mathbb{R}$.
@@
and setting $\xi=\eta$ and integrating produces the trace identity $\sum_{n=0}^{\infty}\lambda_{n}(Q)=4Q$.
@@
For a truncated convolution operator on [Q,Q][-Q,Q] with symbol $\widehat{K}(p)=2\pi\,e^{-|p|}$, the first Szegő limit theorem (see [80, 13]) gives the asymptotic distribution of eigenvalues. Defining the symbol of the Love operator $\mathcal{L}_{Q}=(2\pi)^{-1}\mathcal{K}_{Q}$ as $\sigma(p)=1-e^{-|p|}$, then it follows that the eigenvalues of $\mathcal{L}_{Q}$ are $\mu_{n}=\lambda_{n}/(2\pi)\in(0,1)$.
@@
In the large-$Q$ limit, the operator is locally equivalent to $\mathcal{L}$, and for a self-adjoint operator whose symbol takes values in $[0,1]$, the eigenvalue distribution converges (in the sense of (145)) to the push-forward of the Lebesgue measure on momentum space under the symbol map $p\mapsto e^{-|p|}$. The normalisation factor $1/(2Q)$ is the reciprocal of the interval length, reflecting the density of states per unit length.
@@
since $\log(1/(1-\varepsilon))\sim\varepsilon$ for small $\varepsilon$. This confirms that $\mathcal{O}(Q)$ eigenvalues accumulate at $2\pi$ for large $Q$.
@@
where $\alpha=2[1-\log 2]$ [^5] One evaluates the integral using $\int\arctan(x+R)\,dx=(x+R)\arctan(x+R)-\tfrac{1}{2}\log(1+(x+R)^{2})+C$ and expands for large $R$..
@@
The uncertainty principle for functions supported on [Q,Q][-Q,Q] yields
$\int_{-\infty}^{\infty}|p|^{2}\,|\hat{\tilde{\phi}}_{0}(p)|^{2}\,dp\geq\frac{c}{Q^{2}}\,$, (155)
since $\phi_{0}$ cannot simultaneously be concentrated in both position and frequency. Combined with (154) via Cauchy–Schwarz, this yields $\delta\geq c_{1}/(Q\,(\log Q)^{2})$ after optimising the frequency partition. ∎
@@
with $c_{0}\approx 6.43$ and $c_{1}\approx 0.15$, rather than a pure power law $c/Q^{\alpha}$. An effective exponent $\alpha>1$ may be observed over limited ranges of $Q$ (particularly at small $Q$, where pre-asymptotic effects are significant), but for large $Q$ the data is consistent with the algebraic rate $1/Q$ modulated by a slowly growing logarithmic correction. This logarithmic correction originates from the interplay between the linear zero of the symbol $\sigma(p)=1-e^{-|p|}$ (79) at $p=0$ and the Fisher–Hartwig correction to the Szegő asymptotics (see Section A.3).